I have a problem on my homework that I cannot figure out. The question is "If P=2L+ 2W, L=11 feet, and P=25 feet, what is W?" I've tried to figure it out, but I forgot how to solve an equation that looks like that. Coul you please help?
Answer by mananth(16946)

p=2l+2w
l=11
p=25
plug l & p in the equation
25= 2*11+2w
25=22+2w
add -22 to both sides
25-22=22-22+2w
3=2w
/2
3/2 = w
w= 1.5 ft.